OpenStudy (anonymous):

Find S15 for the series -1 + -3 + -5 + -7 +...

OpenStudy (lyrae):

\[a _{n} = 1 - 2n\]\[\sum_{i=1}^{n} a _{n} = \sum_{i=1}^{15} 1-2i = -225\]